In how many ways can the word SUNDAY be arranged.
How many arrangements begin with S and do not end with Y?

Answer by jrfrunner(365)

assuming there are no duplicated letters, there are n! different ways of ordering a word with n letters
----
the word SUNDAY be arranged in 6!=6*5*4*3*2=720 ways
---
The are 5! ways of ordering the word SUNDAY if the first letter is rstricted to S, because only the 5 remaining ones are allowed to change.
There are 4! ways of ordering the word SUNDAY if the first letter is restricted to S "and" the last letter is restricted to Y
--
Therefore, there are 5!-4! ways that the word SUNDAY can be arranged such that the first letter is S and the last letter "is not" Y
5!-4! = 120-24 =96